# GOFA: A Generative One-For-All Model for Joint Graph Language Modeling

### v1

@@ 6 Conclusion, Limitations, and Future works
We introduce GOFA, a generative One-for-All graph foundation model. GOFA is pre-trained in a graph-level next-token prediction manner to enable large-scale self-supervised learning. By integrating GNN layers with LLM layers, GOFA combines the generative capabilities of LLMs for free-form output with the structural learning strengths of GNNs for understanding complex graph connections. Our experiments demonstrate that GOFA, when fine-tuned with a small number of data, achieves impressive zero-shot performance, highlighting its potential as a robust graph foundation model. One limitation of our work is the extensive training time required due to the use of abundant graph data. Furthermore, interleaving GNN layers with LLM layers increases the embedding dimensions for nodes and edges, which consequently extends the inference time beyond that of a typical LLM. Additionally, we employ a frozen LLM compressor in our architecture; hence, the compression capability is not dynamically integrated with the graph data, potentially impacting the efficiency and adaptability of the model. We believe finetuning a graph language compressor can further enhance the performance of GOFA and will explore it in the future.
@@ A.1 In-context autoencoder (ICAE)
This section briefly introduces ICAE and how it helps build the GOFA model; please refer to ICAE paper Ge et al. [2023] for the specifics of the model. ICAE contains two decoder-only LLMs. One serves as a language compressor that compresses sentences into a fixed-length sequence of vectors, and the other serves as a language decoder that decodes or queries into the compressed sentence representations. Specifically, during training, an input token sequence $\bm{x}=\{x_{1},...,x_{l}\}$ is appended by a $K$ memory tokens $\{m_{1},...,m_{k}\}$ with trainable embeddings. The concatenated sequence is fed to the LLM compressor with a LoRA adapter Hu et al. [2022].
@@
$\begin{split}\{l(m_{1}),...,l(m_{K}),l(x_{1}),...,l(x_{l})\}&=LLM_{dec}(\{h(m_{1}),...,h(m_{K}),e(x_{1}),...,e(x_{l})\})\\ \min_{\Theta_{comp}}\text{CrossEntropy}&(\{l(m_{K}),l(x_{1}),...,l(x_{l-1})\},\{x_{1},...,x_{l}\})\end{split}$ (6)
@@
By training this auto-encoder objective on a large-scale, the compressor model learns to compress all information about a sentence to the memory token outputs like in a conventional auto-encoder model. Because the compressed representation contains as much information as possible, GNN can pass messages between nodes with minimal information loss.
@@ A.3 Transformer Convolutional GNN
As mentioned in Section 3.2, we customize a Transformer Convolutional GNN(TransConv) Shi et al. [2021] as the GNN used in Equation 3. Since GNN layers operate on token representations and tokens at different indices do not communicate, we describe the GNN at one index for simplicity. The $t$-th GNN layer on node $i$ and its neighbors $\mathcal{N}(i)$ is:
@@
$h(\cdot)$ represents input node and edge features. $\bm{W}$ represents query ($q$), key ($k$), value ($v$), output ($o$) linear projection for nodes and edges. The formulation closely follows the transformer design Vaswani et al. [2017] and its GNN adaptation Shi et al. [2021]. This formulation does not aggregate the last layer embedding $h^{t}(i)$ into the next layer, because we already add residual to maintain the same effect. We use pre-layer normalization following Llama Touvron et al. [2023].
Appendix B Supervised Experiment Results
@@
In the supervised experiment, GOFA ’s prompt does not include class optional. We show the supervised results in Table 5. GOFA achieved competitive performance on most datasets. In particular, GOFA achieved SOTA performance on the Pubmed dataset, demonstrating that GOFA can transfer pre-trained knowledge to downstream tasks. The fluidity of GOFA also allows it to perform regression tasks, which is not possible for earlier general graph models like OFA Liu et al. [2023a]. We also notice that GOFA is not performing well on some link datasets, possibly due to the fact that in a supervised setting, we only train a small portion of the data (specific numbers in the experimental details section in Appendix F), and different link tasks require subtle structural information that is difficult to learn and not present in the pre-training tasks. We plan to incorporate more diverse structural tasks into our pre-training datasets for the model to better generalizability.
@@
Table 5: Experiment results in supervised learning.
Cora Cora PubMed PubMed Arxiv WikiCS WN FB Products ML1M
Task type Link Node Link Node Node Node Link Link Node Link
Metric Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ RMSE $\downarrow$
GCN 78.9$\pm$0.6 82.3$\pm$1.1 77.5$\pm$0.4 77.8$\pm$0.7 73.9$\pm$0.6 77.0$\pm$0.6 82.7$\pm$0.4 90.1$\pm$0.3 80.0$\pm$0.7 0.977$\pm$0.001
GAT 80.1$\pm$0.3 80.4$\pm$0.4 80.5$\pm$0.2 76.6$\pm$0.5 75.8$\pm$0.3 79.8$\pm$0.5 88.8$\pm$0.3 93.6$\pm$0.1 81.4$\pm$0.2 0.923$\pm$0.001
OFA 87.97 75.34 95.89 77.89 73.44 77.62 98.31 95.78 - -
GOFA-llama 80.96 73.08 79.80 86.54 77.08 78.92 86.62 81.73 75.96 1.197
GOFA-mistral 83.48 77.08 81.65 87.33 74.28 80.34 88.87 80.11 78.54 1.090
@@ Appendix C Datasets
Cora. The Cora dataset is a co-citation network, where nodes are papers related to artificial intelligence. Edges mean the connected two papers are co-cited by other papers. The Cora dataset contains 2708 nodes and 10556 edges. We collect the Cora dataset and its raw text from OFA Liu et al. [2023a]. We evaluate the performance of the baseline and our proposed model on Cora for both node-level and link-level tasks. For the node-level task, the aim is to classify the node into the correct paper category from 7 different categories. The split is obtained from OFA. It contains 140/500/2068 samples for train/val/test set respectively. For the link-level task, the object is to predict whether two paper nodes are co-cited or not. We follow the setting of OFA Liu et al. [2023a] and randomly split all edges into train/val/test sets with a ratio of 0.85/0.05/0.1.
@@
PubMed. The PubMed dataset is a co-citation network, where nodes are papers related to diabetes mellitus. Edges mean the connected two papers are co-cited by other papers. The PubMed dataset contains 19717 nodes and 88648 edges. We collect the PubMed dataset and its raw text from OFA Liu et al. [2023a]. We evaluate the performance of the baseline and our proposed model on PubMed for both node-level and link-level tasks. For the node-level task, papers have 3 different categories. The goal is to classify the node into the correct paper category. We obtain the split directly from original source. It contains 60/500/19157 samples for train/val/test set respectively. For the link-level task, the object is to predict whether two paper nodes are co-cited or not. We follow the setting of OFA Liu et al. [2023a] and randomly split all edges into train/val/test sets with a ratio of 0.85/0.05/0.1.
@@
Arxiv. The Arxiv dataset is a citation network, where nodes are papers related to computer science and edges mean two papers have a citation relationship. The Arxiv dataset contains 169343 nodes and 1166243 edges. We collect the Arxiv dataset and its raw text from OGB Hu et al. [2021b]. We evaluate the node classification on the Arxiv dataset. The goal is to classify the paper node into the correct category from 40 possible categories. We obtain the split directly from OGB Hu et al. [2021b]. It contains 90941/29799/48603 samples for train/val/test set, respectively.
@@
WikiCS. The WikiCS dataset is a graph obtained from Wikipedia. The nodes in WikiCS are Wikipedia terms and their descriptions. The edges mean there is a hyperlink between two terms. We collect the WikiCS dataset and its raw text from Mernyei and Cangea [2020]. There are 11701 nodes and 216123 edges in the graph. We evaluate the performance of WikiCS on the node classification task. There are 10 different classes. We follow the same split as OFA Liu et al. [2023a], which contains 580/1769/5847 samples for the train/val/test set, respectively.
@@
Products. The Products dataset is a co-purchase graph. The nodes in the graph represent product items from the Amazon platform, and the edges represent that two products are co-purchased together. We obtain the Products and their raw texts from TAPE He et al. [2024a], which is a subset from the original ogbn-Products Hu et al. [2021b] dataset. It contains 54025 nodes and 144638 edges. We evaluate the node classification performance on Products. The data from the original source contains 47 different categories. However, we found that there are two classes that don’t have any nodes and one class contains nodes with missing text. Therefore, we exclude these classes. Finally, there are 44 different categories and 14695/1567/36982 samples for the train/val/test set, respectively.
MoiveLens-1M (ML1M). The ML1M dataset is a bipartite graph obtained from a movie rating platform. The nodes in the graph represent a user or a movie. The edges represent the score the user rates to the movie. The rating ranges from 1 to 5. We obtained the ML1M dataset and its raw texts followed PyG Fey and Lenssen [2019]. It contains 9923 nodes and 1000209 edges. We evaluate the performance of models on ML1M dataset with a link regression task. Specifically, the task is to predict the score the user give to the movie. For ML1M, we randomly split all edges into train/val/test with a ratio 0.85/0.05/0.1, which results in 850177/50011/100021 samples for the train/val/test set, respectively.
WN18RR. The WN18RR is a knowledge graph generated from WordNet. The nodes in the graph represent English words and edges represent the relationship between two words. We obtain the WN18RR dataset from OFA Liu et al. [2023a]. The graph contains 40943 nodes and 93003 relations. We evaluate the performance of baselines and our model on the link classification task using WN18RR. There are 11 different relationships in the dataset. We obtain the split directly from OFA Liu et al. [2023a]. It contains 86835/3034/3134 samples for the train/val/test set, respectively.
@@
FB15K237. The FB15K237 is a knowledge graph generated from Free Base. Nodes in the dataset represent entities in the world and edges represent the relation between entities. We obtained the dataset from OFA Liu et al. [2023a]. The FB15K237 is used to evaluate the link classification. The dataset contains 237 unique classes. We follow the setting of OFA Liu et al. [2023a] and split the dataset with a ratio of 0.85/0.05/0.1, which results in a total of 272115/17535/20466 samples for train/val/test set, respectively.
@@
ExplaGraphs. The ExplaGraphs is a graph question answering dataset on commonsense concepts. Nodes in the dataset represent a common sense concept and edges represent the relation between two concepts. We obtain the dataset from G-retriever He et al. [2024b] The ExplaGraphs can be used for question-answering on graphs. We obtain the split directly from G-retriever He et al. [2024b]. It contains 1659/553/554 graph samples from the train/val/test set.
@@
SceneGraphs. The SceneGraphs is a graph question answering dataset on scene graphs. Nodes in the dataset represent an object in an image and edges represent the relation between two objects. We obtain the dataset from G-retriever He et al. [2024b] The SceneGraphs can be used for question-answering on graphs. We obtain the split directly from G-retriever He et al. [2024b]. It contains 59978/19997/20025 graph samples from the train/val/test set.
@@
MAG240M. The MAG240M dataset is a citation network generated from Microsoft Academic Graphs. The nodes represent academic papers and the links represent a citation relation between two papers. We obtained the dataset and raw text from OGB-lsc Hu et al. [2021a]. However, the original dataset is extremely large and contains nodes without text features (author and institution nodes), since we mainly use the dataset for pre-training, we further downsample the original dataset. Specifically, we only keep paper nodes and citation links between papers. Further, we downsample the edges in the following ways. First, we selected all nodes in the train/val/test split provided by OGB-lsc Hu et al. [2021a]. Next, we filter the edges through two rounds. In the first round, we only keep the edge if either the source or the target is in the selected nodes. If any node in the added edge is not in the selected nodes, we add it to the node set. Next, in the second round, we include additional edges where both the source and target are in the selected nodes (additional nodes are added in the first round). The procedure results in a total of 5875010 nodes and 26434726 edges.
@@
Ultrachat200k. The Ultrachat200k is a question-answering dataset. each sample is a multi-round conversation obtained from the web. We obtained the Ultrachat200k from Ding et al. [2023]. However, the original dataset is not a network. To convert it to a graph dataset, we manually create a graph structure for it. Specifically, if the original sample has $k$ round of conversation, we will generate $k-1$ graph sample. The $i$-th graph will contain the first $i$ round of conversation. Each node in the graph is either a question or an answer. The question and answer are linked by a directed edge indicating the order of the conversation. The conversation of $i+1$ round will be the question-answer pair for this graph. Since we mainly use the dataset for pre-training. We only include train-sft subset. After the conversion, there are a total of 449929 graphs in total.

### v2

@@ 6 Conclusion, Limitations, and Future Works
We introduce GOFA, a generative One-for-All graph foundation model. GOFA is pre-trained under a graph completion framework to enable large-scale self-supervised learning. By integrating GNN layers with LLM layers, GOFA combines the generative capabilities of LLMs for free-form output with the structural learning strengths of GNNs for understanding complex graph connections. Our experiments demonstrate that GOFA, when fine-tuned with a small number of data, achieves impressive zero-shot performance in unseen datasets. However, our zero-shot experiment mainly focuses on evaluating models on unseen datasets with similar task formats instead of unseen task formats. Additionally, we employ a frozen LLM compressor in GOFA ; hence, the compression capability is not naturally unified with the graph data, potentially impacting the adaptability of the model. We will explore harder zero-shot settings and better fine-tuning strategies in the future.
7 Acknowledgement
This work is partially supported by the National Key R&D Program of China (2022ZD0160300) and NSF grant CBE-2225809.
@@ A.1 In-context autoencoder (ICAE)
This section briefly introduces ICAE and how it helps build the GOFA model; please refer to ICAE paper (Ge et al., 2023) for the specifics of the model. ICAE contains two decoder-only LLMs. One serves as a language compressor that compresses sentences into a fixed-length sequence of vectors, and the other serves as a language decoder that decodes or queries into the compressed sentence representations. Specifically, during training, an input token sequence $\bm{x}=\{x_{1},...,x_{l}\}$ is appended by a $K$ memory tokens $\{m_{1},...,m_{k}\}$ with trainable embeddings. The concatenated sequence is fed to the LLM compressor with a LoRA adapter (Hu et al., 2022).
@@
$\begin{split}\{l(m_{1}),...,l(m_{K}),l(x_{1}),...,l(x_{l})\}&=LLM_{dec}(\{h(m_{1}),...,h(m_{K}),e(x_{1}),...,e(x_{l})\}),\\ \min_{\Theta_{comp}}\text{CrossEntropy}&(\{l(m_{K}),l(x_{1}),...,l(x_{l-1})\},\{x_{1},...,x_{l}\}).\end{split}$ (5)
@@
By training this auto-encoder objective on a large-scale, the compressor model learns to compress all information about a sentence to the memory token outputs like in a conventional auto-encoder model. In Table 5, we provide a few examples of the comparison between the original text and the text decoded from the compressed memory tokens by ICAE’s decoder. Because the compressed representation contains as much information as possible, GNN can pass messages between nodes with minimal information loss.
Table 5: Comparison between original texts and decoded text from the compressed representation.
Original Text Decoded Text
Actress Halle Berry has been sharing a number of stunning photos from the time she has spent in Morocco and she just posted a new one to her Instagram page that fans will not want to miss. Halle Berry has been sharing a number of stunning photos from the time she has spent in Morocco and just posted a new one on her Instagram page that fans won’t want to miss.
Utah avoided the turnover bug on Saturday for the first time since its season opener. In addition, the running game was clicking and the defense was dominant as the Utes snapped a two-game winning streak on the road, beating Pittsburgh 26-14. Five keys to Utah’s victory: 1. Utah running back John White IV: Running strong and with purpose from the beginning, White was a big reason why the Utes were within striking distance at halftime. White, who took a couple pops that dislodged his helmet and caused a cut below his ear, seemed to get stronger as the game wore on. He finished the afternoon with 171 yards on 36 carries. Utah avoided the turnover bug on Saturday for the first time since its season opener. In addition, the running game was clicking and the defense was dominant as the Utes snapped a two-game winning streak on the road, beating Pittsburgh 26-14. Five keys to Utah’s victory: 1. 2. 3. 4. 5. Utah running back John John White IV IV ran strong and with purpose from the beginning, being a big reason why the Utes were within striking distance at halftime. He took a couple of shots that dislodged his helmet and caused a cut below his ear, but seemed to get stronger as the game went on. He finished the afternoon with 171 yards on 36 carries.
@@ A.3 Transformer Convolutional GNN
As mentioned in Section 3.2, we customize a Transformer Convolutional GNN(TransConv) (Shi et al., 2021) as the GNN used in Equation 3. Since GNN layers operate on token representations and tokens at different indices do not communicate, we describe the GNN at one index for simplicity. The $t$-th GNN layer on node $i$ and its neighbors $\mathcal{N}(i)$ is:
@@
$h(\cdot)$ represents input node and edge features. $\bm{W}$ represents query ($q$), key ($k$), value ($v$), output ($o$) linear projection for nodes and edges. The formulation closely follows the transformer design (Vaswani et al., 2017) and its GNN adaptation (Shi et al., 2021). This formulation does not aggregate the last layer embedding $h^{t}(i)$ into the next layer, because we already add residual to maintain the same effect. We use pre-layer normalization following Llama (Touvron et al., 2023).
Appendix B Additional Experiments
B.1 Additional zero-shot evaluation results
Table 6: Additional comparison on zero-shot learning results. (The number indicates the number of ways).
Cora-node (7) WikiCS (10) Products (10)
DGI 20.15 13.63 21.35
GraphMAE2 37.08 19.03 25.46
Best-LLM 60.54 63.63 70.16
Best Graph-LLM 69.53 43.45 66.07
GOFA-T 70.81 71.17 79.33
Table 7: Zero-shot learning with molecule dataset.
BBBP HIV
OFA - 35.67
MoMu 49.81 50.26
Galactica 53.94 33.85
GIMLET 59.39 66.24
GOFA 54.91 53.02
To further demonstrate the advantage of GOFA compared to previous methods, we conduct additional zero-shot experiments over different datasets and baselines. First, we additionally compare the zero-shot result with traditional GNN baselines. Since models like GCN and GAT cannot do zero-shot tasks due to their supervised training nature, we compare against graph self-supervised learning baselines. These methods learn node embeddings, which are compared to label text embeddings to make predictions based on cosine similarity. Specifically, we include DGI Veličković et al. (2018) and GraphMAE2 Hou et al. (2023) and follow the same evaluation procedure. The result is shown in Table 7. GOFA significantly outperforms methods that use GNNs for self-supervised learning in the zero-shot setting. This result further demonstrates the effectiveness of the architecture design and the graph modeling tasks of GOFA on zero-shot generalization ability over graph tasks.
Next, we evaluate the GOFA on more numerically and structurally intensive tasks. Specifically, we conduct experiments on molecular property prediction tasks using ogbg-molhiv and BBBP datasets. The experiments closely follow the setting of GIMLET Zhao et al. (2023a). First, we fine-tune the pre-trained GOFA on randomly sampled 100,000 question-answering pairs from the Chembl pretrain dataset. The questions concentrated on asking about different molecule properties. Next, we evaluate the fine-tuned GOFA on ogbg-molhiv and BBBP in a zero-shot setting. We use AUROC as the evaluation metric. For comparison, we include OFA Liu et al. (2023a), MoMu Su et al. (2022), GIMLET Zhao et al. (2023a), and Galactica Taylor et al. (2022). The results are shown in Table 7. We directly report results of OFA Liu et al. (2023a) from their paper, and all other baselines are referred from GIMLET Zhao et al. (2023a). We can see that with only 100,000 instruction-tuning samples, GOFA already achieves comparable results to baselines. Note that GIMLET fine-tunes its model on the whole Chembl-pretrain dataset, which contains more than 400 million question-answering pairs.
B.2 Supervised Experiment Results
@@
In this section, we conduct a supervised learning experiment with the pre-trained GOFA . In the supervised experiment, GOFA ’s prompt does not include class optional. We show the supervised results in Table 8. Specifically, we compare the result of GOFA with the following baselines: 1. basic GNNs, which are trained individually on each dataset, including GCN (Kipf & Welling, 2017) and GAT (Veličković et al., 2018). 2. The contrastive learning methods, including DGI (Veličković et al., 2018) and BGRL (Thakoor et al., 2021). For these methods, we directly report the best result from (He & Hooi, 2024). 3. Graph foundation model, including OFA (Liu et al., 2023a) and UniGraph (He & Hooi, 2024). GOFA achieved competitive performance on most datasets. In particular, GOFA achieved SOTA performance on the Pubmed dataset, demonstrating that GOFA can transfer pre-trained knowledge to downstream tasks. We also notice that GOFA is not performing as well on some datasets, possibly because in a supervised setting, we only train a small portion of the data for one epoch (specific numbers in the experimental details section in Appendix F.5), and in the supervised setting, it is important to see training samples multiple times to ensure detailed understanding of the distribution. As we pre-train the model with more diverse datasets, GOFA can potentially obtain world knowledge as an LLM, which makes transfer learning in the supervised setting more accurate.
@@
Table 8: Experiment results in supervised learning. Bold and underlined shows best and runner-up results.
Cora Cora PubMed PubMed Arxiv WikiCS WN FB Products
Task type Link Node Link Node Node Node Link Link Node
Metric Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$
GCN 78.9$\pm$0.6 82.3$\pm$1.1 77.5$\pm$0.4 77.8$\pm$0.7 73.9$\pm$0.6 77.0$\pm$0.6 82.7$\pm$0.4 90.1$\pm$0.3 80.0$\pm$0.7
GAT 80.1$\pm$0.3 80.4$\pm$0.4 80.5$\pm$0.2 76.6$\pm$0.5 75.8$\pm$0.3 79.8$\pm$0.5 88.8$\pm$0.3 93.6$\pm$0.1 81.4$\pm$0.2
DGI - 51.99$\pm$0.45 - 55.76$\pm$0.56 55.21$\pm$0.21 67.11$\pm$0.12 52.04$\pm$0.22 26.99$\pm$0.22 64.21$\pm$0.32
BGRL - 56.73$\pm$0.23 - 63.77$\pm$0.23 62.21$\pm$0.21 70.12$\pm$0.15 56.44$\pm$0.21 64.91$\pm$0.22 63.77$\pm$0.23
OFA 87.97 75.34 95.89 77.89 73.44 77.62 98.31 95.78 -
UniGraph - 81.43$\pm$0.55 - 74.33$\pm$0.23 72.91$\pm$0.42 79.98$\pm$1.21 85.45$\pm$0.34 94.81$\pm$1.32 80.11$\pm$0.23
GOFA 89.54 76.50 93.97 83.83 74.77 79.96 92.16 88.21 79.98
B.3 Evaluation on additional graph structural tasks
Table 9: Evaluation on additional structural tasks (Acc).
Node degree SPD Node Count
Mistral 98.6 99.8 99.9
GOFA 98.2 99.7 86.3
To further evaluate the potential of GOFA to serve as a general-purpose graph foundational model, we conducted additional experiments on a variety of graph structure tasks, including node degree, shortest path distance, and node count. To provide a fair comparison, we selected Mistral-7B as the baseline model and designed the prompts for Mistral following the incident prompt format proposed in (Fatemi et al., ). Specifically, the graph is described as "[NODE.A] is connected to [NODE.B]". We use Arxiv as the graph dataset and randomly selected 10000 samples for fine-tuning and 2000 samples for testing. For the mistral model, we remove all original text from the dataset, as the model will run out of memory with all the text kept. The evaluation result is shown in Table 9. We can see that GOFA achieved a competitive performance to the LLM baseline in both node-level (node degree) and link-level (shortest path distance) structural tasks. However, LLMs perform better on graph-level tasks. We suspect that the reason lies in the fact that we keep the original text in the graph dataset, which makes the learning of structure much harder, especially when the number of nodes in the graph becomes large. At the same time, our pre-training tasks focus on understanding local graph structures. Thus, GOFA is unfamiliar with understanding global structural information like node count. In the future, we aim to include more structural tasks to further improve the generalization ability of GOFA on different graph tasks.
B.4 Example of GOFA’s Free-form Answer
Figure 6 in the main body illustrates GOFA’s capability to respond to various questions based on the same graph from ogbn-arXiv. In this section, we provide additional examples in Figure 7 to further show the ability of GOFA’s free-form text answer.
Figure 7: Demonstration of GOFA’s ability to respond to any question to the given graph. Above is an example of the products dataset, where the model needs to output the majority category of its connected nodes. Below is another example on Wikics dataset. GOFA is asked to generate a Wikipedia page named based on the graph information.
@@ Appendix C Datasets
Cora. The Cora dataset is a co-citation network, where nodes are papers related to artificial intelligence. Edges mean the connected two papers are co-cited by other papers. The Cora dataset contains 2708 nodes and 10556 edges. We collect the Cora dataset and its raw text from OFA (Liu et al., 2023a). We evaluate the performance of the baseline and our proposed model on Cora for both node-level and link-level tasks. For the node-level task, the aim is to classify the node into the correct paper category from 7 different categories. The split is obtained from OFA. It contains 140/500/2068 samples for train/val/test set respectively. For the link-level task, the object is to predict whether two paper nodes are co-cited or not. We follow the setting of OFA (Liu et al., 2023a) and randomly split all edges into train/val/test sets with a ratio of 0.85/0.05/0.1.
@@
PubMed. The PubMed dataset is a co-citation network, where nodes are papers related to diabetes mellitus. Edges mean the connected two papers are co-cited by other papers. The PubMed dataset contains 19717 nodes and 88648 edges. We collect the PubMed dataset and its raw text from OFA (Liu et al., 2023a). We evaluate the performance of the baseline and our proposed model on PubMed for both node-level and link-level tasks. For the node-level task, papers have 3 different categories. The goal is to classify the node into the correct paper category. We obtain the split directly from original source. It contains 60/500/19157 samples for train/val/test set respectively. For the link-level task, the object is to predict whether two paper nodes are co-cited or not. We follow the setting of OFA (Liu et al., 2023a) and randomly split all edges into train/val/test sets with a ratio of 0.85/0.05/0.1.
@@
Arxiv. The Arxiv dataset is a citation network, where nodes are papers related to computer science and edges mean two papers have a citation relationship. The Arxiv dataset contains 169343 nodes and 1166243 edges. We collect the Arxiv dataset and its raw text from OGB (Hu et al., 2021b). We evaluate the node classification on the Arxiv dataset. The goal is to classify the paper node into the correct category from 40 possible categories. We obtain the split directly from OGB (Hu et al., 2021b). It contains 90941/29799/48603 samples for train/val/test set, respectively.
@@
WikiCS. The WikiCS dataset is a graph obtained from Wikipedia. The nodes in WikiCS are Wikipedia terms and their descriptions. The edges mean there is a hyperlink between two terms. We collect the WikiCS dataset and its raw text from (Mernyei & Cangea, 2020). There are 11701 nodes and 216123 edges in the graph. We evaluate the performance of WikiCS on the node classification task. There are 10 different classes. We follow the same split as OFA (Liu et al., 2023a), which contains 580/1769/5847 samples for the train/val/test set, respectively.
@@
Products. The Products dataset is a co-purchase graph. The nodes in the graph represent product items from the Amazon platform, and the edges represent that two products are co-purchased together. We obtain the Products and their raw texts from TAPE (He et al., 2024a), which is a subset from the original ogbn-Products (Hu et al., 2021b) dataset. It contains 54025 nodes and 144638 edges. We evaluate the node classification performance on Products. The data from the original source contains 47 different categories. However, we found that there are two classes with missing labels. To be consistent with previous literature, we adopt the approach in LLaGA to replace the label name with special symbols. There are 14708/1572/37745 samples for the train/val/test set, respectively.
@@
FB15K237. The FB15K237 is a knowledge graph generated from Free Base. Nodes in the dataset represent entities in the world and edges represent the relation between entities. We obtained the dataset from OFA (Liu et al., 2023a). The FB15K237 is used to evaluate the link classification. The dataset contains 237 unique classes. We follow the setting of OFA (Liu et al., 2023a) and split the dataset with a ratio of 0.85/0.05/0.1, which results in a total of 272115/17535/20466 samples for train/val/test set, respectively.
@@
ExplaGraphs. The ExplaGraphs is a graph question answering dataset on commonsense concepts. Nodes in the dataset represent a common sense concept and edges represent the relation between two concepts. We obtain the dataset from G-retriever (He et al., 2024b) The ExplaGraphs can be used for question-answering on graphs. We obtain the split directly from G-retriever (He et al., 2024b). It contains 1659/553/554 graph samples from the train/val/test set.
@@
SceneGraphs. The SceneGraphs is a graph question answering dataset on scene graphs. Nodes in the dataset represent an object in an image and edges represent the relation between two objects. We obtain the dataset from G-retriever (He et al., 2024b) The SceneGraphs can be used for question-answering on graphs. We obtain the split directly from G-retriever (He et al., 2024b). It contains 59978/19997/20025 graph samples from the train/val/test set.
@@
MAG240M. The MAG240M dataset is a citation network generated from Microsoft Academic Graphs. The nodes represent academic papers and the links represent a citation relation between two papers. We obtained the dataset and raw text from OGB-lsc (Hu et al., 2021a). However, the original dataset is extremely large and contains nodes without text features (author and institution nodes), since we mainly use the dataset for pre-training, we further downsample the original dataset. Specifically, we only keep paper nodes and citation links between papers. Further, we downsample the edges in the following ways. First, we selected all nodes in the train/val/test split provided by OGB-lsc (Hu et al., 2021a). Next, we filter the edges through two rounds. In the first round, we only keep the edge if either the source or the target is in the selected nodes. If any node in the added edge is not in the selected nodes, we add it to the node set. Next, in the second round, we include additional edges where both the source and target are in the selected nodes (additional nodes are added in the first round). The procedure results in a total of 5875010 nodes and 26434726 edges.
@@
Ultrachat200k. The Ultrachat200k is a question-answering dataset. each sample is a multi-round conversation obtained from the web. We obtained the Ultrachat200k from (Ding et al., 2023). However, the original dataset is not a network. To convert it to a graph dataset, we manually create a graph structure for it. Specifically, if the original sample has $k$ round of conversation, we will generate $k-1$ graph sample. The $i$-th graph will contain the first $i$ round of conversation. Each node in the graph is either a question or an answer. The question and answer are linked by a directed edge indicating the order of the conversation. The conversation of $i+1$ round will be the question-answer pair for this graph. Since we mainly use the dataset for pre-training. We only include train-sft subset. After the conversion, there are a total of 449929 graphs in total.
Wikikg90m. Wikikg90m is an encyclopedic knowledge graph dataset extracted from Wikidata knowledge base. We obtain the original Wikikg90m from OGB-LSC (Hu et al., 2021a). It contains 91,230,610 entities, 1,387 relations, and 601,062,811 edges.
WikiGraph. The WikiGraph dataset is designed to increase the diversity of the training texts. Hence, we use WikiText (Merity et al., 2022) dataset as the seed dataset. It contains plain sentences from Wikipedia pages. We generate graphs for sentences with more than 500 characters. Specifically, we first prompt an LLM to extract meaningful entities or concepts from a sentence, and these entities become the nodes in the graph. We then randomly pair concepts to generate edges. Again, we use LLM to generate a description of the relationship between the paired concepts and use the description as the edge text.
BBBP: The BBBP is a molecule property prediction dataset. The task is to predict the effectiveness of molecule to brain blood barrier. We collect the dataset from GIMLET Zhao et al. (2023a). The dataset contains 1631/204/204 molecule graphs for the train/validation/test set, respectively. We directly use the test set for zero-shot evaluation.
HIV: The HIV dataset is a molecule property prediction dataset. The task is to predict the effectiveness of molecules to hiv. We collect the dataset from GIMLET Zhao et al. (2023a). The dataset contains 32901/4113/4113 molecule graphs for the train/validation/test set, respectively. We directly use the test set for zero-shot evaluation.
Chembl pretrain: The Chembl pretrain dataset is a large-scale molecule property prediction dataset. for each simple molecule, the dataset contains several questions to ask about the different properties of the molecule. Overall, it contains 400 million question-answering pairs. We collected it from GIMLET Zhao et al. (2023a) and use it for fine-tuning GOFA for molecule zero-shot experiments.